OpenStudy (anonymous):

Rework problem 1 in section 4.1 of your text, involving the flipping of a loaded coin, but assume that Pr[H] = 0.3. Also, assume that the coin is flipped 4 times, and the random variable X is defined to be 3 times the number of heads minus 2 times the number of tails. How many different values are possible for the random variable X?

OpenStudy (perl):

can you wait, brb

OpenStudy (perl):

lets make an auxilary variable. Let Z = # of heads flipped in 4 coins. so Z can equal 0,1,2,3,4 . then X = 3*Z - 2*(4-Z)

OpenStudy (perl):

because if Z is the number of heads , then 4-Z is the number of tails.

OpenStudy (perl):

like zero heads = 4 tails ,

OpenStudy (perl):

so simplifying X = 3Z - 8 + 2Z X = 5Z - 8 so you have X = -8, -3, 2, 7, 12